The maximum volume a weather balloon can reach without rupturing is 22 000L. It is designed to reach an altitude of 35 km. At this altitude, the atmospheric pressure is 0.0125 atm. The atmospheric pressure at the launch site is 98.85 kPa.
Calculate the maximum volume of helium that the weather balloon should be filled with before it is launched.
1
Expert's answer
2020-06-23T02:16:18-0400

Use Boyle's Law

"P_1V_1=P_2V_2"

convert atm to kPa

"0.0125 atm \times (\frac{101.325 kPa}{1 atm})=1.267 kPa"


"1.267 kPa\times 22000L = 98.85 kPa\times V_2"

"V_2 = 282.0 L"